OpenStudy (anonymous):

(1/8x^2-1/4x-3/8)+(-1/2x^2-1/4x-1/2) Add the following polynomials

OpenStudy (anonymous):

(1/8x^2-1/4x-3/8)+(-1/2x^2-1/4x-1/2) \(\large \frac{1}{8}x^2-\frac{1}{4}x-\frac{3}{8} \) + \(\large -\frac{1}{2}x^2-\frac{1}{4}x-\frac{1}{2} \) _________________________________________________ \(\large (\frac{1}{8}+-\frac{1}{2})x^2+(-\frac{1}{4}+-\frac{1}{4})x+(-\frac{3}{8}+-\frac{1}{2}) \) can you simplify from here?